# An Adaptive Online Smoother with Closed-Form Solutions and Information-Theoretic Lag Selection for Conditional Gaussian Nonlinear Systems

### v1

@@ 4.3 Online Parameter Estimation of Nonlinear Systems with Only Partial Observations
Parameter estimation and model identification for complex nonlinear dynamical systems are crucial for effective state estimation, DA, and forecasting. In partially observed nonlinear systems, parameter and state estimation typically occur simultaneously, for which their solution often requires complicated approximations and expensive numerical methods. Nevertheless, the closed analytic formulas of the online nonlinear smoother (3.11)–(3.12) facilitate an efficient online parameter estimation scheme for the CGNS (2.1)–(2.2) for when only the time series of $\mathbf{x}$ is observed. In this section, we incorporate the adaptive online smoother into an expectation-maximization (EM) algorithm [133, 134, 135, 136, 137] to alternately estimate parameters and hidden states, employing closed analytic formulas for both estimations.
@@ 4.3.1 The Online EM Parameter Estimation Algorithm
The basic EM algorithm is outlined in Appendix C, and the mathematical details for the CGNS can be found in [138]. This reference serves as the foundation for developing an online version that utilizes the adaptive-lag online smoother for simultaneous state estimation and model identification, and also includes the tools necessary for incorporating into this algorithm’s skeleton extensions like block decomposition for learning high dimensional systems, sparse identification or regularization, and physical and other constraints on the parameters.
@@
$\hat{\boldsymbol{\theta}}_{\text{MLE}}:=\underset{\boldsymbol{\theta}\in\Theta}{\mathrm{argmax}}\left\{\log\left(p\big{(}\mathbf{x};\boldsymbol{\theta}\big{)}\right)\right\}=\underset{\boldsymbol{\theta}\in\Theta}{\mathrm{argmax}}\left\{\log\left(\int_{\mathbf{y}}p\big{(}\mathbf{x},\mathbf{y};\boldsymbol{\theta}\big{)}{\,\rm d}\mathbf{y}\right)\right\},$ (4.6)
where $\Theta$ denotes the underlying parameter space, which we assume to be a closed and convex subset of $\mathbb{C}^{N}$ for $N\in\mathbb{N}$, and the integration in (4.6), i.e., marginalization, is happening over the state space of $\mathbf{y}$. The EM iteration alternates between performing the so-called expectation step (E-step), which estimates the hidden state of $\mathbf{y}$ using the current estimate for the parameters, and a maximization step (M-step), which computes the parameters that maximize the expected marginal log-likelihood found in the E-step [135, 139, 140]. Since $\mathbf{y}$ is unobserved throughout time, the conditional expectation in the E-step needs to be taken for those measurable functions (measurable with respect to the current observable $\sigma$-algebra) containing $\mathbf{y}$ when computing the log-likelihood estimate, which takes into account the uncertainty in the state estimation through the optimal a-posteriori statistics of $\mathbf{y}$. As is shown in Appendix C, the solution from the E-step is exactly the online smoother posterior distribution.
@@ 4.3.2 A Numerical Example
${\,\rm d}u(t)$ $=(-d_{u}u(t)+\gamma u(t)v(t)+F_{u}){\,\rm d}t+\sigma_{u}{\,\rm d}W_{u}(t),$ (4.7)
${\,\rm d}v(t)$ $=(-d_{v}v(t)-\gamma u(t)^{2}+F_{v}){\,\rm d}t+\sigma_{v}{\,\rm d}W_{v}(t),$ (4.8)
where $d_{u},d_{v},\gamma_{1},\gamma_{2},F_{u},F_{v}\in\mathbb{R}$ and $\sigma_{u},\sigma_{v}>0$, with $W_{u}$ and $W_{v}$ being two mutually independent Wiener processes. The system (4.7)–(4.8) fits the CGNS framework for $\mathbf{x}:=u$ and $\mathbf{y}=v$. Similar to the model in (4.1)–(4.2) of Section 4.1, the variable $v$ acts as a stochastic damping in the observable process, which leads to intermittent extreme events. The effect of these instabilities in the convergence of the algorithm is especially studied. We assume a time series of $u$ is continuously observed while there is no observation of $v$. To simplify the study in this section, we assume that the noise feedbacks $\sigma_{u}$ and $\sigma_{v}$ are known, although our framework does not require such a restriction [138]. Thus, the parameter vector for this model is as follows:
$\boldsymbol{\theta}:=(d_{u},\gamma_{1},F_{u},d_{v},\gamma_{2},F_{v})^{\mathtt{T}}.$ (4.9)
@@
$d_{v}=1,\quad F_{v}=0.8,\quad\sigma_{v}=1.$
@@
The numerical setup is as follows. The initial guess of the parameter values is $\boldsymbol{\theta}_{0}=(2,6,2,0.5,6,0.4)^{\mathtt{T}}$. The numerical integration time step, which is also the observational frequency in such a continuous observational case, is $\Delta t=0.001$. The total length of observations is $T=200$ time units. Since implementing the online smoother requires a certain history of previous state estimation solutions, the observed time series of $u$ within a short initial window $[0,T_{\text{ini}}]$, with $T_{\text{ini}}=10$, is utilized to initially carry out the filtering solution and online smoother estimates, based on $\boldsymbol{\theta}_{0}$. This can be treated as a burn-in or learning period for the algorithm, and also aids in avoiding the possibility of short-time blowup. Afterwards, with the arrival of each new observation, we update the adaptive-lag online smoother state estimations via the procedure in (3.27)–(3.28) and then re-estimate the parameters using the MLE. The update happens at every observation as to ensure stability. The hyperparameters in the adaptive-lag method are $b=1000$ (where $b\Delta t=1$ time unit) and $\delta=10^{-4}$, where we use the actual sequence of information gains to define the adaptive lag at each new observation, i.e., (3.33).
@@
Figure 4.6 presents the results from using the online EM algorithm. Panels (a) and (b) display the true trajectories of the observed and unobserved variables, $u$ and $v$, respectively, displayed in blue. These trajectories are generated by running equations (4.7)–(4.8) forward in time. Alongside the true trajectories, the time series obtained by simulating the model with the estimated parameters at the final iteration of the online EM algorithm are included in red. Note that the same values for the Wiener processes are used for these two signals, allowing us to have a point-wise comparison. The recovery of the signal for both observable and unobserved variables is notably skillful. It is worth highlighting that, in panel (b), the dashed horizontal lines indicate the true (in black) and online EM estimated (in green) anti-damping threshold lines, above which the unobserved variable effectively acts as anti-damping in the observable process, initiating intermittent instabilities in the signal. Their consistency is critical for retrieving the true dynamics using the estimated parameters. Panel (c) shows the adaptive lag values, measured in time units (i.e., $L_{n}\Delta t$) for the online smoother, illustrating that longer lags are typically required during the onset of the intermittent phases as to effectively capture the dynamics. The efficacious recovery of the true state is also demonstrated through higher-order statistics, such as the PDFs and ACFs for both $u$ (in panels (d)–(e)) and $v$ (in panels (f)–(g)). Finally, panels (h)–(m) present the trace plots produced by the online EM iterations for various parameters of interest. The parameters in the observed process $u$ are nearly perfectly recovered, including the nonlinearity feedback $\gamma$ and the deterministic constant forcing $F_{v}$ in the process of the unobserved variable. The only parameter that is not perfectly recovered is the damping $d_{u}$ in the process of the unobserved variable, with the algorithm converging to twice its true value. This is expected, as these parameters, along with the noise feedback, are typically the hardest to estimate due to observability issues, where their contribution to the observed process is relatively weak. Nevertheless, this inaccuracy does not hinder the model with the estimated parameters from reproducing key dynamical and statistical features, as shown in panels (a)–(b) and (d)–(g).
@@
An important observation from the trace plots is that extreme events in the observable signal enhance the convergence skill of the online EM algorithm within the CGNS framework. In panel (a), four extreme events are marked as A, B, C, and D, characterized by a high signal-to-noise ratio. The trace plots in panels (h)–(m) demonstrate that extreme events influence parameter convergence. The high signal-to-noise ratio mitigates the uncertainty in observations and results in strong observability of the system, which naturally facilitates parameter estimation [141]. This is particularly evident in the anti-damping threshold (panel (m)), highlighting the role of these events in identifying the intermittent nature of the underlying dynamics.
@@
Figure 4.6: Panel (a): True trajectory of the observable variable $u$ over the simulation period (in blue), and the recovered trajectory from using the same white noise values used to generate the true signal (by utilizing the same seed number) but where the model in (4.7)–(4.8) is instead defined through the recovered parameter values at the last iteration of the online EM algorithm (in red). The dashed rectangular box denotes the period of running the filter and online smoother algorithms with $\boldsymbol{\theta}_{0}$, which is $[0,10]$. Four extreme events are marked in the observed time series (denoted by A, B, C, and D). Panel (b): Same as (a) but for the unobserved variable $v$. The true and online EM estimated anti-damping threshold $d_{u}/\gamma_{1}$ are marked by dashed lines, in black and green, respectively. Panel (c): Adaptive lag values generated using (3.33) (measured in time units, i.e., $L_{n}\Delta t$), as a function of simulation time. Panels (d)–(g): PDFs of $u$ (panel (d)) and $v$ (panel (f)) and their ACFs (panels (e) and (g), respectively), for both the true model parameters and the ones obtained at the last iteration of the online EM algorithm. Panels (h)–(m): Trace plots produced by the online EM algorithm.
@@ 5 Conclusions
In this paper, a computationally efficient algorithm for online adaptive-lag optimal smoother-based state estimation with partial observations is developed. Closed analytical formulae are available for this online smoother. It applies to CGNS, representing a rich class of CTNDSs with wide applications in neuroscience, ecology, atmospheric science, geophysics, and many other fields. The adaptive-lag strategy allows a significant reduction of computational storage. Notably, an information-theoretic criterion, exploiting uncertainty reduction across the entire posterior distribution, is developed to calculate the adaptive lag value, which helps in preserving dynamical and statistical consistency.
@@
The adaptive-lag online smoother has been applied to systems with intermittent instabilities (Section 4.1). It effectively recovers extreme events and discovers the causal dependence between different state variables. It is also applicable to high-dimensional systems, such as the LDA problem. The state estimation based on such an online smoother shows nearly the same accuracy as the standard forward-backward smoother while requiring significantly less computational storage (Section 4.2). In addition, the online adaptive-lag smoother facilitates real-time parameter estimation. The study highlights the importance of utilizing observed extreme events to accelerate the convergence of parameter estimation (Section 4.3).
@@
Potential future research topics in this direction are as follows. The closed-form expressions of the online smoother estimates can aid in assessing extreme events and their impact on the CGNS dynamics, as well as in identifying the most sensitive directions that exhibit rapid fluctuations during specific events using information-theoretic methods [98]. This is facilitated by the fact that the CGNS framework enjoys analytic formulae for such conditional distributions [36]. Second, the adaptive-lag online smoother can serve as a useful tool for causal inference and information flow analysis. On the one hand, one can exploit intermittent extreme events to improve state estimation, which helps reduce the model error in developing parsimonious surrogate models via causal inference. On the other hand, the conditional Gaussianity of the framework allows the application of transfer or causation entropy to explore these topics in future studies. Finally, future work also involves developing a rigorous investigation of the spectral properties of the update tensors $\mathbf{D}^{j,n-2}$ used in the online smoother, as well as of the matrices $\mathbf{E}^{j}$ that define them, potentially through suitable concentration inequalities. This includes studying the temporal behavior of the information-theoretic criterion for adaptive lags, $\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)$ (and its LSDF). Numerical case studies have shown that these quantities increase exponentially over time, with the rate depending on whether a quiescent or intermittent signal is observed. Notably, the spectrum of $\mathbf{D}^{j,n-2}$ remains below the convergence threshold of $1$, which is reassuring. A rigorous analysis of these properties is a natural future topic.

### v2

@@ 4.3 Online Parameter Estimation of Nonlinear Systems with Only Partial Observations
Parameter estimation and model identification for complex nonlinear dynamical systems are crucial for effective state estimation, DA, and forecasting. In partially observed nonlinear systems, parameter and state estimation typically occur simultaneously, requiring a solution that often necessitates complicated approximations and expensive numerical methods. Nevertheless, the closed analytic formulae of the online nonlinear smoother (3.10)–(3.11) facilitate an efficient online parameter estimation scheme for the CGNS (2.1)–(2.2) for when only the time series of $\mathbf{x}$ is observed. In this section, we incorporate the adaptive online smoother into an expectation-maximization (EM) algorithm [103, 104, 105, 106] to alternately estimate parameters and hidden states, employing closed analytic formulae for both estimations.
@@ 4.3.1 The Online EM Parameter Estimation Algorithm
The basic EM algorithm is outlined in Appendix E, and the mathematical details for the CGNS framework can be found in [107]. This reference serves as the foundation for developing an online version that utilizes the adaptive-lag online smoother for simultaneous state estimation and model identification, and also includes the tools necessary for incorporating into this algorithm’s skeleton extensions like block decomposition for learning high-dimensional systems, sparse identification or regularization, and physical and other constraints on the parameters.
@@
$\hat{\boldsymbol{\theta}}_{\text{MLE}}:=\underset{\boldsymbol{\theta}\in\Theta}{\mathrm{argmax}}\left\{\log\big(p(\mathbf{x};\boldsymbol{\theta})\big)\right\}=\underset{\boldsymbol{\theta}\in\Theta}{\mathrm{argmax}}\left\{\log\left(\int_{\mathbf{y}}p\big(\mathbf{x},\mathbf{y};\boldsymbol{\theta}\big){\,\rm d}\mathbf{y}\right)\right\},$ (4.7)
where $\Theta$ denotes the underlying parameter space, which we assume to be a closed and convex subset of $\mathbb{C}^{N}$ for $N\in\mathbb{N}$, and the integration in (4.7), i.e., marginalization, is happening over the state space of $\mathbf{y}$. The EM iteration alternates between performing the so-called expectation step (E-step), which estimates the hidden state of $\mathbf{y}$ using the current estimate for the parameters, and a maximization step (M-step), which computes the parameters that maximize the expected marginal log-likelihood found in the E-step [104, 108, 109]. Since $\mathbf{y}$ is unobserved throughout time, the conditional expectation in the E-step needs to be taken for those measurable functions (measurable with respect to the current observable $\sigma$-algebra) containing $\mathbf{y}$ when computing the log-likelihood estimate, which takes into account the uncertainty in the state estimation through the optimal a-posteriori statistics of $\mathbf{y}$. As is shown in Appendix E, the solution from the E-step is exactly the online smoother posterior distribution.
@@ 4.3.2 A Numerical Example
${\,\rm d}u(t)$ $=\big(-d_{u}u(t)+\gamma u(t)v(t)+F_{u}\big){\,\rm d}t+\sigma_{u}{\,\rm d}W_{u}(t),$ (4.8)
${\,\rm d}v(t)$ $=\big(-d_{v}v(t)-\gamma u(t)^{2}+F_{v}\big){\,\rm d}t+\sigma_{v}{\,\rm d}W_{v}(t),$ (4.9)
where $d_{u},d_{v},\sigma_{u},\sigma_{v}>0$ and $\gamma,F_{u},F_{v}\in\mathbb{R}$, with $W_{u}$ and $W_{v}$ being two mutually independent Wiener processes. The system (4.8)–(4.9) fits the CGNS framework for $\mathbf{x}:=u$ and $\mathbf{y}=v$. Similar to the model in (4.1)–(4.2) of Section 4.1, the variable $v$ acts as a stochastic damping in the observable process, which leads to intermittent extreme events. The effect of these instabilities in the convergence of the algorithm is especially studied. We assume a time series of $u$ is continuously observed while there is no observation of $v$. To simplify the study in this section, we assume that the noise feedbacks $\sigma_{u}$ and $\sigma_{v}$ are known, although our framework does not require such a restriction [107]. Thus, the parameter vector for this model is as follows:
$\boldsymbol{\theta}:=(d_{u},\gamma,F_{u},d_{v},F_{v})^{\mathtt{T}}.$ (4.10)
@@
$d_{v}=1,\quad F_{v}=0.2,\quad\sigma_{v}=1.$
@@
The numerical setup is as follows. The initial guess of the parameter values is $\boldsymbol{\theta}_{0}=(2,6,2,0.5,0.6)^{\mathtt{T}}$. The numerical integration time step, which is also the observational frequency in such a continuous observational case, is $\Delta t=0.001$. The total length of observations is $T=200$ time units. Since implementing the online smoother requires a certain history of previous state estimation solutions, the observed time series of $u$ within a short initial window $0,T_{\text{ini}}]$, with $T_{\text{ini}}=10$, is utilized to initially carry out the filtering solution and online smoother estimates, based on $\boldsymbol{\theta}_{0}$. This can be treated as a burn-in or learning period for the algorithm, and also aids in avoiding the possibility of short-time blowup. Afterwards, with the arrival of each new observation, we update the adaptive-lag online smoother state estimations via the procedure in (3.26)–(3.27) and then re-estimate the parameters using the MLE. The update happens at every observation to ensure stability. The hyperparameters in the adaptive-lag method are $b=1000$ (where $b\Delta t=1$ time unit) and $\delta=10^{-4}$, where we use the actual sequence of information gains to define the adaptive lag at each new observation, i.e., (3.32). For a brief discussion on the online EM parameter estimation algorithm’s stability, sensitivity, and convergence skill with respect to the initial parameter values and length of the burn-in or learning period (for this specific model), through a purely empirical and simplified analysis, the interested reader is referred to Appendix F.
@@
Figure 4.6 presents the results from using the online EM algorithm. Panels (a) and (b) display the true trajectories of the observed and unobserved variables, $u$ and $v$, respectively, displayed in blue. These trajectories are generated by running equations (4.8)–(4.9) forward in time. Alongside the true trajectories, the time series obtained by simulating the model with the estimated parameters at the final iteration of the online EM algorithm are included in red. Note that the same values for the Wiener processes are used for these two signals, allowing us to have a point-wise comparison. The recovery of the signal for both observable and unobserved variables is notably skillful. It is worth highlighting that, in Panel (b), the dashed horizontal lines indicate the true (in black) and online EM estimated (in green) anti-damping threshold lines, above which the unobserved variable effectively acts as anti-damping in the observable process, initiating intermittent instabilities in the signal. Their consistency is critical for retrieving the true dynamics using the estimated parameters. Panel (c) shows the adaptive lag values, measured in time units (i.e., $L_{n}\Delta t$) for the online smoother, illustrating that longer lags are typically required during the onset of the intermittent phases to effectively capture the dynamics. The efficacious recovery of the true state is also demonstrated through higher-order statistics, such as the PDFs and autocorrelation functions (ACFs) for both $u$ (in Panels (d)–(e)) and $v$ (in Panels (f)–(g)). The use of an upper lag bound of $b\Delta t=1$ time unit is enough to capture most of the information in the online smoother updates while not overspending on storage (for the chosen tolerance value of $\delta=10^{-4}$). This is showcased in the produced PDFs and ACFs, which indicate faithful model fidelity and memory recovery, respectively, where for the latter, this is true even for time-lags outside the interval of the produced adaptive lag values. Finally, Panels (h)–(m) present the trace plots produced by the online EM iterations for various parameters of interest. The parameters in the observed process $u$ are nearly perfectly recovered, including the nonlinear feedback $\gamma$ and the deterministic constant forcing $F_{v}$ in the process of the unobserved variable. The only parameter that is not perfectly recovered is the damping $d_{v}$ in the process of the unobserved variable, with the algorithm converging to twice its true value. This is expected, as these parameters, along with the noise feedback, are typically the hardest to estimate due to observability issues, where their contribution to the observed process is relatively weak. Nevertheless, this inaccuracy does not hinder the model with the estimated parameters from reproducing key dynamical and statistical features, as shown in Panels (a)–(b) and (d)–(g).
@@
An important observation from the trace plots is that extreme events in the observable signal enhance the convergence skill of the online EM algorithm within the CGNS framework. In Panel (a), four extreme events are marked as A, B, C, and D, characterized by a high signal-to-noise ratio. The trace plots in Panels (h)–(m) demonstrate that extreme events influence parameter convergence. The high signal-to-noise ratio mitigates the uncertainty in observations, resulting in strong observability of the system, which naturally facilitates parameter estimation [110]. This is particularly evident in the anti-damping threshold (Panel (m)), highlighting the role of these events in identifying the intermittent nature of the underlying dynamics.
@@
Figure 4.6: Panel (a): True trajectory of the observable variable $u$ over the simulation period (in blue), and the recovered trajectory from using the same white noise values used to generate the true signal (by utilizing the same seed number) but where the model in (4.8)–(4.9) is instead defined through the recovered parameter values at the last iteration of the online EM algorithm (in red). The dashed rectangular box denotes the period of running the filter and online smoother algorithms with $\boldsymbol{\theta}_{0}$, which is $[0,10]$. Four extreme events are marked in the observed time series (denoted by A, B, C, and D). Panel (b): Same as (a) but for the unobserved variable $v$. The true and online EM estimated anti-damping threshold $d_{u}/\gamma$ are marked by dashed lines, in black and green, respectively. Panel (c): Adaptive lag values generated using (3.32) (measured in time units, i.e., $L_{n}\Delta t$), as a function of simulation time. Panels (d)–(g): PDFs of $u$ (Panel (d)) and $v$ (Panel (f)) and their ACFs (Panels (e) and (g), respectively), for both the true model parameters and the ones obtained at the last iteration of the online EM algorithm. Panels (h)–(m): Trace plots produced by the online EM algorithm.
For the sake of completeness, in Figure G.1, we provide the results from instead using the fixed-lag online discrete smoother in the outlined online EM parameter estimation algorithm, when applied to this specific model identification problem. See Appendix G for more info and a discussion comparing the fixed- and adaptive-lag online EM parameter estimation results.
@@ 5 Conclusions
In this paper, a computationally efficient algorithm for online adaptive-lag optimal smoother-based state estimation with partial observations is developed. Closed analytical formulae are available for this online smoother. It applies to CGNS, representing a rich class of CTNDSs with wide applications in neuroscience, ecology, atmospheric science, geophysics, and many other fields. Importantly, the adaptive-lag strategy allows the reduction of computational complexity and storage requirements, creating the potential for the algorithm to significantly outperform both offline smoothing and its fixed-lag counterpart operationally while achieving a comparable skill in state estimation. Notably, an information-theoretic criterion, exploiting uncertainty reduction across the entire posterior distribution, is developed to calculate the adaptive lag value, which helps in preserving dynamical and statistical consistency.
@@
The adaptive-lag online smoother has been applied to a system with intermittent instabilities (Section 4.1). It effectively recovers extreme events and discovers the causal dependence between different state variables. It is also applicable to high-dimensional systems, such as the LDA problem. The state estimation based on such an online smoother shows nearly the same accuracy as the standard forward-backward smoother while requiring significantly less computational storage under suitable choices of the algorithm parameters (Section 4.2). In addition, the online adaptive-lag smoother facilitates real-time parameter estimation. The study highlights the importance of utilizing observed extreme events to accelerate the convergence of parameter estimation (Section 4.3).
@@
Potential future research topics in this direction are as follows. First, the closed-form expressions of the online smoother estimates can aid in assessing extreme events and their impact on the CGNS dynamics, as well as in identifying the most sensitive directions that exhibit rapid fluctuations during specific events using information-theoretic methods [78]. This is facilitated by the fact that the CGNS framework enjoys analytic formulae for such conditional distributions [33]. Second, the adaptive-lag online smoother can serve as a useful tool for causal inference and information flow analysis. On the one hand, one can exploit intermittent extreme events to improve state estimation, which helps reduce the model error in developing parsimonious surrogate models via causal inference. On the other hand, the conditional Gaussianity of the framework allows the application of transfer or causation entropy to explore these topics in future studies. Third, it is necessary to explicitly investigate how the adaptive-lag online smoother has the potential to outperform the fixed-lag alternative by uniformly reducing storage needs and computational complexity in CGNSs. Such a study will involve the derivation of theoretical or empirical representations of these operational metrics, as well as of point-wise or distribution-based error measures during state estimation, as functions of the fixed lag and adaptive-lag tolerance parameter $\delta$ (for a fixed upper lag bound $b$). Finally, future work along this line also involves developing a rigorous investigation of the spectral properties of the update tensors $\mathbf{D}^{j,n-2}$ used in the online smoother, as well as of the matrices $\mathbf{E}^{j}$ that define them, potentially through suitable concentration inequalities. This includes studying the temporal behavior of the information-theoretic criterion for adaptive lags, $\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)$ (and its LSDev). Numerical case studies have shown that these quantities showcase exponentially varying behavior over time, with the rate depending on whether a quiescent or intermittent signal is observed. Notably, the spectrum of $\mathbf{D}^{j,n-2}$ remains below the convergence threshold of $1$, which is reassuring. A rigorous analysis of these properties is a natural future topic.
@@ Acknowledgments
The authors thank the reviewers for their constructive feedback and valuable suggestions during the peer-review process.